An 15 g ice cube at -23˚C is put into a Thermos flask containing 150 cm3 of water at 17˚C. By how much has the entropy of the cube-water system changed when a final equilibrium state is reached? The specific heat of ice is 2200 J/kg K and that of liquid water is 4187 J/kg K. The heat of fusion of water is 333 × 103 J/kg.

Please show explanation and answer.
Answer #1

initially the ice warms from -23oC to 0oC
The change in entropy in this case is
       ΔS1 = integral ( dQ /T)
              = m Cice ln (Tf / Ti)
              = (15 x 10-3) (2200) ln (273 / 250)
              = 2.90 J/ K
Now the ice melts at 0oC , Then
       ΔS2 = Q / T = mL /T   = 15 * 333 / 273 = 18.3 J /K
Now the water warms to lake temperature
         ΔS3 = m Cw ln(Tf / Ti)
                = (15 / 1000) (4187) ln (290 / 273)
                = 3.793 J /K
Therefore total change in entropy
         ΔS = 2.90 +18.3 + 3.793 = 25 J /K